A boy's outfit consists of slacks, shirt, and shoes. He does not want to wear the same combination on any two days of the year. Presuming that he has the same number of slacks, shirts, and pairs of shoes, what is the minimum number of each that he will need?
(a) 3 (b) 6 (c) 7 (d) 8 (e) 9


Answer by richwmiller(17219)

s*s*s=365
s^3=365
s=7.1
s > 7
s=8
7^3=343
8^3=512